# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 2 Algorithmic Framework and Setup for Analysis
Consider the variant of the algorithm where we run the cluster-based rounding procedure with probability $\frac{\alpha}{2}$, and the pivot-based procedure with threshold $1/3$ with the remaining probability of $1-\frac{\alpha}{2}$. By Lemma 7, the expected cost of the clustering given by the variant is at most
$\quad\sum_{uv\in E^{+}}\left(\frac{\alpha}{2}\cdot\frac{2x_{uv}}{1+x_{uv}}+\big{(}1-\frac{\alpha}{2}\big{)}\cdot b^{+}_{\alpha}(x_{uv})\right)+\sum_{uv\in E^{-}}\left(\frac{\alpha}{2}\cdot\frac{1-x_{uv}}{1+x_{uv}}+\big{(}1-\frac{\alpha}{2}\big{)}\cdot b^{-}_{\alpha}(x_{uv})\right)$
@@
As a baseline, we provide a per-triangle analysis leading to an approximation ratio of $1.5$ in Section 5:
@@
For budget functions $b^{+}\equiv b^{+}_{1.5}$ and $b^{-}\equiv b^{-}_{1.5}$, we have $\mathrm{cost}(T)\leq\Delta(T)$ for every triangle $T$.
Clearly, the lemma implies that (4) holds for $b^{+}\equiv b^{+}_{1.5}$ and $b^{-}\equiv b^{-}_{1.5}$. By Lemma 8, our algorithm gives an approximation ratio of $1.5$. We remark that $1.5$ is the best possible ratio we can achieve using the per-triangle analysis. For a $+$\!+\!-+ + - triangle with length $\frac{1}{2}$ for $+$edges and length $1$ for the --edge, we need to pay a factor of $2$ for each of the $\frac{1}{2}$-length $+$edge. Then the cluster-based rounding algorithm gives factors of $2$ and $\frac{4}{3}$ for $+$edges of lengths 00 and $\frac{1}{2}$ respectively. For the pivot-based rounding algorithm, the factors are at least 00 and $2$. A combination of the two algorithms can only lead to a factor of $1.5$.
To get a better approximation ratio, we provide two analyses that use global distributions of triangles. The former is purely analytic and the latter relies on solving a factor-revealing SDP. The following two lemmas are proved in Sections 6 and 7 respectively.
@@
(4) holds for budget functions $b^{+}\equiv b^{+}_{1.49}$ and $b^{-}\equiv b^{-}_{1.49}$.
Lemma 11.
(4) holds for budget functions $b^{+}\equiv b^{+}_{1.437}$ and $b^{-}\equiv b^{-}_{1.437}$.
Combined with Lemma 8, the two lemmas imply Theorems 2 and 3 respectively.
3 Overview of Techniques
In this section, we provide overviews of the techniques used in our results.
Simpler and Better Preclustering Procedure.
The concept of preclustering was introduced in [CLLN23]. In a preclustered instance, we predetermine the fate of some edges: for some edges $uv$, $u$ and $v$ must be in the same cluster; for some other edges $uv$, $u$ and $v$ must be separated. Since the relation of being in the same cluster is transitive, we define a preclustered instance using a pair $({\mathcal{K}},E_{\mathrm{adm}})$, where ${\mathcal{K}}$ is a partition of $V$ into so called atoms and $E_{\mathrm{adm}}\subseteq{V\choose 2}$ is a set of admissible edges. An atom can not be broken. If $u$ and $v$ are not in the same atom and $uv\notin E_{\mathrm{adm}}$, then $u$ and $v$ must be separated. [CLLN23] showed how to construct a preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, losing only a $(1+\varepsilon)$ factor in the optimum cost, while at the same time guaranteeing that $|E_{\mathrm{adm}}|\leq O({\mathrm{opt}}/\varepsilon^{12})$. This is crucial for their correlated rounding algorithm, as it loses an additive error depending on $|E_{\mathrm{adm}}|$. In this work, we still need the preclustering procedure to bound the rounding error, but now it is inside the procedure of solving the cluster LP.
We greatly simplify the preclustering procedure from [CLLN23], and as a result, we achieve a much better bound of $O({\mathrm{opt}}/\varepsilon^{2})$ on $|E_{\mathrm{adm}}|$. [CLLN23] used the agreement graph to construct the atoms; roughly speaking, two vertices are in agreement if their neighborhood sets are similar to each other. The analysis uses many technical structural lemmas from [CLM+21], which solves Correlation Clustering in the online setting. In contrast, our construction of atoms is simple: we construct an $O(1)$-approximate clustering ${\mathcal{C}}$, mark vertices whose costs are large, and then ${\mathcal{K}}$ is obtained from ${\mathcal{C}}$ by removing marked vertices and creating singletons for them. The set of admissible edges is roughly defined as follows: we construct a graph $(V,E^{1})$ where two vertices are neighbors if their $+$degrees are similar. Then an edge $uv$ is admissible if $u$ and $v$ have many common neighbors in $E^{+}\cap E^{1}$.
Solving Cluster LP by Preclustering.
As we mentioned, we move the complication of handling rounding errors to the step of solving the cluster LP. As in [CLLN23], we construct a preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$, and formulate an LP relaxation aimed at finding the $(1+\varepsilon)$-approximate good clustering for $({\mathcal{K}},E_{\mathrm{adm}})$, that we call the bounded sub-cluster LP. In contrast to [CLLN23], which solves many instances of this LP embedded in their round-or-cut framework, we only solve the LP once, therefore avoiding this heavy framework. With a solution $(x,y)$ to the LP, we run a procedure that constructs a single cluster $C$ randomly. The probability that any vertex is in $C$ is precisely $1/y_{\emptyset}$, where $y_{\emptyset}$ is the fractional number of clusters in $y$. The probabilities that exactly one of $u$ and $v$ is in $C$, and both of them are in $C$, are respectively $\frac{x_{uv}}{y_{\emptyset}}$ and $\frac{1-x_{uv}}{y_{\emptyset}}$ up to some error terms arising from the Raghavendra-Tan rounding procedure. As usual, $x_{uv}$ is the extent in which $u$ and $v$ are separated.
To construct the solution $z=(z_{S})_{S\subseteq V}$ for the cluster LP, we generate $y_{\emptyset}\Delta$ many clusters $C$ independently, for a large enough polynomial $\Delta$. Roughly speaking, the solution $z$ is $\frac{1}{\Delta}$ times the multi-set of clusters $C$ we generated. The error incurred by the Raghavendra-Tan rounding procedure can be bounded in terms of $|E_{\mathrm{adm}}|$, and the error from sampling can be bounded using concentration bounds.
1.49-approximation.
We start with the algorithm of [CLLN23], but make several key modifications both in the design and in the analysis. This allows us to significantly improve the approximation ratio, first to 1.5 and, eventually, to 1.49, which shows that, perhaps surprisingly, even the rather low approximation factor of $1.5$ is not tight for Correlation Clustering. The first key ingredient is to use a principled budget function for the pivot-based rounding procedure, defined earlier in (5), which is designed to optimally balance the approximation factor of edges between the two rounding procedures. This new budget function is better than the one used in [CLLN23], but does not allow us to reach 1.5 without changing the algorithm. Indeed, the budget for the short $+$edges in ++++\!+\!++ + + triangles is still too low to reach the approximation ratio 1.5. Thus, the second key ingredient is to add the threshold step to the pivot-based rounding procedure for the short $+$edges (i.e., $+$edges $uv$ with $x_{uv}\leq 1/3$). By adding this threshold step, the cost of the triangles containing such edges decreases; for example, a ++++\!+\!++ + + triangle with all short edges now has cost zero. This allows us to use the new budget function and still reach 1.5. Notice that making the threshold too large would result in too much cost for $+$\!+\!-+ + - triangles.
Finally, we observe that, analogous to the correlated rounding approach of [CLN22], only the bad triangles are tight, meaning their cost equals their budget. Roughly speaking, a bad triangle is a $+$\!+\!-+ + - triangle whose two $+$edges have value very close to half and whose --edge has value close to one. This allows us to apply a charging argument, in which tight triangles have part of their cost paid for by triangles that are not tight (i.e., that have extra budget). Now there are no tight triangles (i.e., all triangles have some unused budget), and we can decrease the $\alpha$ in the budget function from $1.5$ to $70/47$. As previously [ACN08, CMSY15, CLN22, CLLN23], the analysis necessary to reach 1.5 and go below requires a case-by-case analysis of triangle types to ensure that the budget allocated to each triangle covers its cost. Both the new threshold step and the new budget functions result in an analysis that is more involved than what was required in [CLLN23], but is still feasible.
1.437-approximation.
The above charging argument between different types of triangles can be more systematically expressed by a factor-revealing SDP. Given a cluster LP solution $z_{S}$ and vertices $u,v,w$, let us define $y_{uv}:=\sum_{S\supseteq\{u,v\}}z_{S}$ (resp. $y_{uvw}:=\sum_{S\supseteq\{u,v,w\}}z_{S})$ be the probability that $u,v$ (resp. $u,v,w$) are in the same cluster. Given any quadruple $T=(a,b,c,d)\in[0,1]^{4}$, let $\eta_{T}$ represent the number of triangles $(u,v,w)$ such that of $y_{uv}=a,y_{uw}=b,y_{vw}=c,y_{uvw}=d$. The above 1.49-approximation analysis can be regarded as putting one constraint on the distribution of $\eta_{T}$. To enhance the approximation ratio and reduce the budget function, we opt for a more detailed categorization of triangles, imposing stronger constraints on $\eta_{T}$.
@@ 3 Overview of Techniques
Consider an imaginary rounding procedure, where given a pivot $u$, the cluster $C$ that contains $u$ is simply chosen with probability $z_{C}$ (note that $\sum_{C\ni u}z_{C}=1$). Let $X_{v}$ denote the event that node $v$ is included in the cluster of node $u$ in this rounding. We can show $\operatorname*{{\mathbb{E}}}[X_{v}\cdot X_{w}]=y_{uvw}$ and $\operatorname*{{\mathbb{E}}}[X_{v}]\cdot\operatorname*{{\mathbb{E}}}[X_{w}]=y_{uv}y_{uw}$. The covariance matrix $COV_{u}$, where $COV_{u}(v,w)=\operatorname*{{\mathbb{E}}}[X_{v}\cdot X_{w}]-\operatorname*{{\mathbb{E}}}[X_{v}]\cdot\operatorname*{{\mathbb{E}}}[X_{w}]=y_{uvw}-y_{uv}y_{uw}$, must be positive semidefinite (PSD). This PSD constraint on the covariance matrix enforces a stronger constraint on $\eta_{T}$. For instance, if all non-degenerate triangles centered at $u$ are $+$\!+\!-+ + - triangles with $y$ value $(y_{uv}=0.5,y_{uw}=0.5,y_{wv}=0,y_{uvw}=0)$, then the covariance matrix of $COV_{u}$ cannot be PSD because $COV_{u}(v,w)=y_{uvw}-y_{uv}y_{uw}=-0.25$ for almost all non-diagonal entries.
@@
Despite there being infinitely many types of triangles in each range $I_{i},I_{j},I_{k}$, our key observation is that $y_{uvw}-y_{uv}y_{uw}$ is multi-linear. Therefore, we only need a few triangles in each range to represent all possible triangles. We want to mention the triangles we need are fixed so can be precomputed and the only unsure variable is $\eta_{T}$. To compute a lower bound $\sum\eta_{T}(\Delta(T)-\text{cost}(T))$, we set up a semi-definite program (SDP) under the constraint that $Q$ is PSD. This SDP is independent of cluster LP and relies on the chosen interval and budget function. By employing a practical SDP solver, we demonstrate that $\sum\eta_{T}(\Delta(T)-\text{cost}(T))\geq 0$.
Gaps and Hardness.
A high-level intuition for the cluster LP is the following: (any) LPs cannot distinguish between a random graph and a nearly bipartite graph. For the cluster LP, given a complete graph $H=(V_{H},E_{H})$ with $n=|V_{H}|$, our Correlation Clustering instance is $G=(V_{G},E_{G})$ where $V_{G}=E_{H}$ and $e,f\in V_{G}$ have a plus edge in $G$ if they share a vertex in $V$. Consider vertices of $H$ as ideal clusters in $G$ containing their incident edges. The LP fractionally will think that it is nearly bipartite, implying that the entire $E_{H}$ can be partitioned into $n/2$ ideal clusters of the same size. Of course, integrally, such a partition is not possible in complete graphs.
@@
Still, we use a similar construction for $3$-uniform hypergraphs; let $H=(V_{H},E_{H})$ be the underlying $3$-uniform hypergraph and $G=(V_{G},E_{G})$ be the plus graph of the final Correlation Clustering instance where $V_{G}=E_{H}$ and $e,f\in E_{H}$ has an edge in $G$ if they share a vertex in $H$. We use the hardness result of Cohen-Addad, Karthik, and Lee [CAL22] that shows that it is hard to distinguish whether $H$ is nearly bipartite, which implies that half of the vertices intersect every hyperedge, or close to a random hypergraph.
@@
We show how to solve the cluster LP in Section 4, proving Theorem 1. It uses our improved preclustering procedure, which will be described in Section A. We prove Lemmas 9, 10 and 11 in Sections 5, 6 and 7 respectively; they will prove Theorems 2 and 3. We give the $(\frac{4}{3}-\varepsilon)$-integrality gap of the cluster LP (Theorem 4) in Section 8, and the improved hardness of $24/23-\varepsilon$ (Theorem 5) in Section 9.
@@
For two sets $A$ and $B$, we use $A\triangle B=(A\setminus B)\cup(B\setminus A)$ to denote the symmetric difference between $A$ and $B$. We used $N^{+}_{u}$ and $N^{-}_{u}$ to denote the sets of $+$ and --neighbors of a vertex $u$ respectively in the Correlation Clustering instance. For a clustering ${\mathcal{C}}$ of $V$, we define $\mathrm{obj}({\mathcal{C}})$ to be the objective value of ${\mathcal{C}}$. For any $x\in[0,1]^{{V\choose 2}}$, we already defined $\mathrm{obj}(x)=\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$. Recall that we defined $\mathrm{cost}_{u}(v,w),\Delta_{u}(v,w),\mathrm{cost}(T)$ and $\Delta(T)$ for a triangle $T=(u,v,w)$ or a degenerate triangle $T=(u,v)$ in Section 2; they depend on the budget functions $b^{+}$ and $b^{-}$.
4 Solving Cluster LP Relaxation Approximately
4.1 Preclustering
We use the definition of a preclustered instance from [CLLN23], with some minor modifications.

### v3

@@ 2.2 Rounding Cluster LP
Consider the variant of the algorithm where we run the cluster-based rounding procedure with probability $\frac{\alpha}{2}$, and the pivot-based procedure with threshold $\tau=0.4$ with the remaining probability of $1-\frac{\alpha}{2}$. By Lemma 7, the expected cost of the clustering given by the variant is at most
$\quad\sum_{uv\in E^{+}}\left(\frac{\alpha}{2}\cdot\frac{2x_{uv}}{1+x_{uv}}+\big(1-\frac{\alpha}{2}\big)\cdot\mathrm{budget}^{+}_{\alpha}(x_{uv})\right)+\sum_{uv\in E^{-}}\left(\frac{\alpha}{2}\cdot\frac{1-x_{uv}}{1+x_{uv}}+\big(1-\frac{\alpha}{2}\big)\cdot\mathrm{budget}^{-}_{\alpha}(x_{uv})\right)$
@@
Algorithm 3 Pivot-Based Rounding with Threshold $\tau=0.4$ and $I^{+}_{c}=(\tau,1]$
[…]
2:while $V^{\prime}\neq\emptyset$ do
3:  pick a pivot $u\in V^{\prime}$ uniformly at random;  $C\leftarrow\{u\}$
4:  sample a set $S\ni u$ according to probabilities $\{z_{S}\}$ $\triangleright$ $\sum_{S\ni u}z_{S}=1$
5:  for each $v\in V^{\prime}\cap N^{+}(u)$ do
6:   if $x_{uv}\leq\tau$ then $\triangleright$ independent rounding on out-of-range ++ edges
7:     independently add $v$ to $C$ $\triangleright$ $f^{+}(x)=1$
8:   else if $v\in S$ then $\triangleright$ dependent rounding on in-range ++ edges
9:     add $v$ to $C$
10:  for each $v\in V^{\prime}\cap N^{-}(u)$ do
11:   independently add $v$ to $C$ with probability $1-x_{uv})$ $\triangleright$ $f^{-}(x)=x$
12:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{C\}$;  $V^{\prime}\leftarrow V^{\prime}\setminus C$
13:return $\mathcal{C}$}
As a baseline, we instantiate Algorithm 2 to obtain Algorithm 3. In this instantiation we set $I_{c}^{+}=(0.4,1]$: consequently, all ++-edges $(u,v)$ with $x_{uv}\leq 0.4$ are rounded independently, while ++-edges with $x_{uv}\in I_{c}^{+}$ participate in the dependent rounding step. (All --edges are rounded independently.)
We provide a per-triangle analysis leading to an approximation ratio of $1.56$ in Section 5:
@@
For Algorithm 3 and budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{1.56}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{1.56}$, we have $\mathrm{cost}(T)\leq\Delta(T)$ for every triangle $T$.
Clearly, the lemma implies that (4) holds for $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{1.56}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{1.56}$. By Lemma 8, Algorithm 3 gives an approximation ratio of $1.56$.
To get a better approximation ratio, we provide an analysis that uses global distributions of triangles. It relies on solving a factor-revealing SDP. The resulting instantiation is given in Algorithm 4. For a ++ edge $(u,v)$, we proceed as follows: (i) if $x_{uv}\leq 0.40$, include $v$ in the pivot cluster deterministically; (ii) if $0.40<x_{uv}\leq 0.57$, include $v$ via the dependent step using the pivot-centered set distribution $\{z_{S}\}$; (iii) if $x_{uv}>0.57$, include $v$ independently with probability $1-x_{uv}$. For a - edge $(u,v)$, we split the endpoints with probability $x_{uv}^{2}$ (equivalently, add $v$ to the pivot’s cluster with probability $1-x_{uv}^{2}$). For algorithm 4, the following lemma is proved in Section 6.
Algorithm 4 Pivot-Based Rounding (with correlated mid-range for ++ edges)
[…]
2:while $V^{\prime}\neq\emptyset$ do
3:  pick a pivot $u\in V^{\prime}$ uniformly at random;  $C\leftarrow\{u\}$
4:  sample a set $S\ni u$ according to probabilities $\{z_{S}\}$ $\triangleright$ $\sum_{S\ni u}z_{S}=1$
5:  for each $v\in V^{\prime}\cap N^{+}(u)$ do
6:   if $x_{uv}\leq 0.4$ then $\triangleright$ independent rounding on out-of-range ++ edges
7:     independently add $v$ to $C$ $\triangleright$ $f^{+}(x)=1$ for $x\leq 0.4$
8:   else if $x_{uv}>0.57$ then $\triangleright$ independent rounding on out-of-range ++ edges
9:     add $v$ to $C$ with probability $1-x_{uv}$ $\triangleright$ $f^{+}(x)=x$ for $x>0.57$
10:   else if $v\in S$ then $\triangleright$ dependent rounding on in-range ++ edges
11:     add $v$ to $C$
12:  for each $v\in V^{\prime}\cap N^{-}(u)$ do
13:   independently add $v$ to $C$ with probability $1-x_{uv}^{2}$ $\triangleright$ $f^{-}(x)=x^{2}$
14:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{C\}$;  $V^{\prime}\leftarrow V^{\prime}\setminus C$
15:return $\mathcal{C}$
@@
For Algorithm 4, (4) holds for budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{1.485}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{1.485}$.
Combined with Lemma 8, it implies Theorem 3.
Global Triangle Distributions.
The idea of global triangle distributions, which is a basis for Lemma 10, is systematically expressed by a factor-revealing SDP. Given a cluster LP solution $z_{S}$ and vertices $u,v,w$, let us define $y_{uv}:=\sum_{S\supseteq\{u,v\}}z_{S}$ (resp. $y_{uvw}:=\sum_{S\supseteq\{u,v,w\}}z_{S})$ be the probability that $u,v$ (resp. $u,v,w$) are in the same cluster. Given any quadruple $T=(a,b,c,d)\in[0,1]^{4}$, let $\eta_{T}$ represent the number of triangles $(u,v,w)$ such that of $y_{uv}=a,y_{uw}=b,y_{vw}=c,y_{uvw}=d$.
@@
Consider a hypothetical rounding procedure, where given a pivot $u$, the cluster $C$ that contains $u$ is simply chosen with probability $z_{C}$ (note that $\sum_{C\ni u}z_{C}=1$). Let $X_{v}$ denote the event that node $v$ is included in the cluster of node $u$ in this rounding. We can show $\operatorname*{{\mathbb{E}}}[X_{v}\cdot X_{w}]=y_{uvw}$ and $\operatorname*{{\mathbb{E}}}[X_{v}]\cdot\operatorname*{{\mathbb{E}}}[X_{w}]=y_{uv}y_{uw}$. The covariance matrix $COV_{u}$, where $COV_{u}(v,w)=\operatorname*{{\mathbb{E}}}[X_{v}\cdot X_{w}]-\operatorname*{{\mathbb{E}}}[X_{v}]\cdot\operatorname*{{\mathbb{E}}}[X_{w}]=y_{uvw}-y_{uv}y_{uw}$, must be positive semidefinite (PSD). This PSD constraint on the covariance matrix enforces a stronger constraint on $\eta_{T}$. For instance, if all non-degenerate triangles centered at $u$ are +++\!+\!- triangles with $y$ value $(y_{uv}=0.5,y_{uw}=0.5,y_{wv}=0,y_{uvw}=0)$, then the covariance matrix of $COV_{u}$ cannot be PSD because $COV_{u}(v,w)=y_{uvw}-y_{uv}y_{uw}=-0.25$ for almost all non-diagonal entries.
@@
Despite there being infinitely many types of triangles in each range $I_{i},I_{j},I_{k}$, our key observation is that $y_{uvw}-y_{uv}y_{uw}$ is multilinear. Therefore, we only need a few triangles in each range to represent all possible triangles. We want to mention the triangles we need are fixed so they can be precomputed and the only unsure variable is $\eta_{T}$. To compute a lower bound $\sum\eta_{T}(\Delta(T)-\text{cost}(T))$, we set up a semidefinite program (SDP) under the constraint that $Q$ is PSD. This SDP is independent of cluster LP and relies on the chosen interval and budget function. By employing a practical SDP solver, we demonstrate that $\sum\eta_{T}(\Delta(T)-\text{cost}(T))\geq 0$.
2.3 Gaps and Hardness
A high-level intuition for the cluster LP is the following: (any) LP cannot distinguish between a random graph and a nearly bipartite graph. For the cluster LP, given a complete graph $H=(V_{H},E_{H})$ with $n=|V_{H}|$, our Correlation Clustering instance is $G=(V_{G},E_{G})$ where $V_{G}=E_{H}$ and $e,f\in V_{G}$ are connected by a ++edge in $G$ if they share a vertex in $V$. Consider vertices of $H$ as ideal clusters in $G$ containing their incident edges. The LP fractionally will think that it is nearly bipartite, implying that the entire $E_{H}$ can be partitioned into $n/2$ ideal clusters of the same size. Of course, integrally, such a partition is not possible in complete graphs.
@@ 2.3 Gaps and Hardness
Still, we use a similar construction for $3$-uniform hypergraphs; let $H=(V_{H},E_{H})$ be the underlying $3$-uniform hypergraph and $G=(V_{G},E_{G})$ be the plus graph of the final Correlation Clustering instance where $V_{G}=E_{H}$ and $e,f\in E_{H}$ has an edge in $G$ if they share a vertex in $H$. We use the hardness result of Cohen-Addad, Karthik, and Lee [CKL22] that shows that it is hard to distinguish whether $H$ is nearly bipartite, which implies that half of the vertices intersect every hyperedge, or close to a random hypergraph.
@@
In Section 3, we describe our preclustering procedure. It will be used in our algorithm to solve the cluster LP described in Section 4. This finishes the proof of Theorem 1. We prove Lemma 10 in Section 6, which proves Theorem 3. It is a computer-assisted proof, and we present an analytic proof of Lemma 9, giving a bound of $1.56$, in Section 5. We give the $(\frac{4}{3}-\varepsilon)$-integrality gap of the cluster LP (Theorem 4) in Section 7, and the improved hardness of $24/23-\varepsilon$ (Theorem 5) in Section 8.
@@
For two sets $A$ and $B$, we use $A\triangle B=(A\setminus B)\cup(B\setminus A)$ to denote the symmetric difference between $A$ and $B$. We used $N^{+}_{u}$ and $N^{-}_{u}$ to denote the sets of ++ and -neighbors of a vertex $u$ respectively in the Correlation Clustering instance. For a clustering $\mathcal{C}$} of $V$, we define $\mathrm{obj}({\mathcal{C}})$ to be the objective value of $\mathcal{C}$}. For any $x\in[0,1]^{{V\choose 2}}$, we already defined $\mathrm{obj}(x)=\sum_{uv\in E^{+}}x_{uv}+\sum_{uv\in E^{-}}(1-x_{uv})$. Recall that we defined $\mathrm{cost}_{u}(v,w),\Delta_{u}(v,w),\mathrm{cost}(T)$ and $\Delta(T)$ for a triangle $T=(u,v,w)$ or a degenerate triangle $T=(u,v)$ in this section; they depend on the budget functions $\mathrm{budget}^{+}$ and $\mathrm{budget}^{-}$.
3 Preclustering
An important technique we use to solve the cluster LP is preclustering. We define a preclustered instance and a good clustering as follows.